n 2016, the automobile industry announced the average vehicle selling price was $28,819. Six years earlier, the average price was $22,125.

What was the annual percentage increase in vehicle selling price? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Annual percentage increase in selling price            
Answer #1

Solution:

Let annual percentage increase in selling price = i

Period = 6 years

Average price of vehicle before 6 years =$22,125

Average price of vehicle today = $28,819

$22,125 * (1+i)^6 = $28,819

(1+i)^6 = 1.30255367

1+i = (1.30255367)^1/6 = 1.0450

i = 4.50%